# Phase-space negativity as a computational resource for quantum kernel methods

### v1

@@ Abstract
Quantum kernel methods are a proposal for achieving quantum computational advantage in machine learning. They are based on a hybrid classical-quantum computation where a function called the quantum kernel is estimated by a quantum device while the rest of the computation is performed classically. Quantum advantages may be achieved through this method only if the quantum kernel function cannot be estimated efficiently on a classical computer. In this paper, we provide sufficient conditions for the efficient classical estimation of quantum kernel functions for bosonic systems. Specifically, we show that if the negativity in the phase-space quasi-probability distributions of data-encoding quantum states associated with the quantum kernel scales at most polynomially with the size of the quantum circuit, then the kernel function can be estimated efficiently classically. We consider quantum optical examples involving linear-optical networks with and without adaptive non-Gaussian measurements and investigate the effects of loss on the efficiency of the classical simulation. Our results underpin the role of the negativity in phase-space quasi-probability distributions as an essential resource in quantum machine learning based on kernel methods.
I Introduction
@@
One approach to address this challenge, especially for quantum sampling problems, has been based on the use of quasi-probability distributions [11, 12, 13, 14, 15, 16]. In this approach, a classical description of a quantum experiment is possible as long as one can find probabilistic descriptions for the input state, the evolution, and the measurement in terms of non-negative quasi-probability distributions. Also, the relation between the negativity of quasi-probability distributions and the computational overhead in estimating an output probability of a quantum circuit using those distributions has been investigated [15, 16].
@@
In this paper, we build upon these results to quantify the sampling complexity in estimating quantum kernels in terms of the negativity of the quasi-probability approach. We present sufficient conditions for efficient classical estimation of kernel functions in continuous-variable quantum machine learning. Informally, we show that one can achieve the same performance as a quantum circuit by using classical Monte Carlo-type sampling techniques if the negativity in the phase-space quasi-probability distributions (PQD) of data-encoding quantum states scales at most polynomially with the size of the circuit. This condition is independent of the measurement strategy and holds even if sampling from the output probability distribution of the circuit is classically hard, such as in boson sampling problems.
@@
In light of this, we identify the negativity of PQDs as a necessary resource to achieve quantum speedups in machine learning based on kernel methods. As illustrating examples, we discuss efficient classical estimation of quantum kernel functions that are based on Gaussian states, non-Gaussian output states of linear optical networks, and partially measured Gaussian states, including adaptive measurement strategies. For all these examples, we obtain fine-grained characterizations of the relevant quantum computational resources based on existing non-classicality measures, namely negativity volume [17] and non-classical depth [18, 19].
@@
Our results include a new approach to classical estimation of quantum kernel functions that takes into account the structure of state preparation. A surprising consequence is that estimating quantum kernel functions based on a large class of partially measured Gaussian states can be done efficiently by classical computers despite these states being highly non-Gaussian (see Theorem 1).
@@
The structure of the paper is as follows: in Section II, we set the stage and provide some background on quantum kernel methods; in Section III, we derive general expressions for quantum kernel functions for bosonic systems based on PQDs (see Lemma 1); in Section IV, we introduce classical Monte Carlo sampling methods and derive two phase-space-inspired classical algorithms for estimating quantum kernel functions (see Algorithms 1 and 2), together with rigorous performance guarantees; we illustrate the flexibility and applicability of our algorithms in Section V, by applying them to several examples of high significance in bosonic quantum information processing, identifying the relevant quantum computational resources in each setting; we conclude in Section VI.
II Quantum Kernel Methods
@@
III Kernel functions in phase space
@@
where $\bm{\alpha}$ is a row vector of $m$ complex numbers, where $\text{d}^{2m}\!\bm{\bm{\alpha}}\,=\text{d}^{2m}\!\bm{\,\Re(\bm{\alpha})}\,\text{d}^{2m}\!\bm{\,\Im(\bm{\alpha})}\,$, and where the $\bm{s}$-ordered phase-space quasi-probability distributions [($\bm{s}$)-PQDs] are defined by
@@
Here, $\bm{\alpha},\bm{\xi}$ are row vectors of $m$ complex numbers, the diagonal matrix $\bm{s}=\mathrm{diag}(s_{1},s_{2},\ldots,s_{m})$ specifies the ordering for each mode, and $D(\bm{\xi})=\exp(\bm{\xi}\bm{a}^{\dagger}-\bm{a}\bm{\xi}^{\dagger})$ is the $m$-mode displacement operator with $\bm{a}=(a_{1},\ldots,a_{m})$ and $\bm{a}^{\dagger}=(a_{1}^{\dagger},\ldots,a_{m}^{\dagger})^{T}$ being vectors of annihilation and creation operators, respectively. Note that these formulas are valid for more general ordering matrices, but we restrict to diagonal ones for simplicity.
@@
$\mathrm{Tr}[\rho_{\Pi(x)}\rho_{\Pi(x^{\prime})}]=\pi^{m+2k}\!\!\int_{\bm{\alpha},\bm{\beta}\in\mathbb{C}^{k}}\!\!\!\!\!\!\!\!\!\!\mathrm{d}^{2k}\bm{\alpha}\,\mathrm{d}^{2k}\bm{\beta}\,W_{\Pi(x)}^{(-\bm{u})}(\bm{\alpha})W_{\Pi(x^{\prime})}^{(\bm{v})}(\bm{\beta})\int_{\bm{\gamma}\in\mathbb{C}^{m}}\!\!\!\!\!\!\!\mathrm{d}^{2m}\bm{\gamma}\,W_{U(x)\rho_{\rm{in}}(x)U(x)^{\dagger}}^{(\bm{u}\oplus\bm{s})}(\bm{\alpha},\bm{\gamma})W_{U(x^{\prime})\rho_{\rm{in}}(x^{\prime})U(x^{\prime})^{\dagger}}^{(-\bm{v}\oplus-\bm{s})}(\bm{\beta},\bm{\gamma}).$ (13)
@@
$=\mathrm{Tr}[\sigma(\Pi\otimes\Delta^{(\bm{s})}(\bm{\gamma}))]$ (15)
$=\pi^{k}\!\!\int_{\bm{\alpha}\in\mathbb{C}^{k}}\!\!\!\!\!\mathrm{d}^{2k}\bm{\alpha}W_{\Pi}^{(-\bm{u})}\!(\bm{\alpha})\mathrm{Tr}[\sigma(\Delta^{(\bm{u})}\!(\bm{\alpha})\!\otimes\!\Delta^{(\bm{s})}\!(\bm{\gamma}))]$
@@
IV Classical estimation of kernel functions
In this section, we describe two classical algorithms using Monte Carlo sampling methods to estimate kernel functions based on the expressions obtained in Lemma 1.
IV.1 Monte Carlo estimation
@@
where $\mathcal{N}(W^{(\bm{t})}_{\rho}):=\int_{\mathbb{C}^{n}}\text{d}^{2n}\!\bm{\mu}\,\,\big{|}W^{(\bm{t})}_{\rho}(\bm{\mu})\big{|}$ is a measure of negativity in the ($\bm{t}$)-PQD. We note that $\mathcal{N}(W^{(\bm{t})}_{\rho})=1$ if the distribution is non-negative. Next, using the quasi-probability distribution $W^{(-\bm{t})}_{A}(\bm{\mu})$ representing operator $A$, we introduce the estimator
@@
then with probability at least $1-\delta$ the classical estimate of the expectation value has additive error less than $\epsilon$. Here, the range of the estimator is given by
$\mathcal{R}(E)=\mathcal{N}\big{(}W^{(\bm{t})}_{\rho}\big{)}\mathcal{R}(W^{(-\bm{t})}_{A}),$ (21)
with $\mathcal{R}(W^{(-\bm{t})}_{A})=\pi^{n}[\max_{\bm{\mu}}W^{(-\bm{t})}_{A}(\bm{\mu})-\min_{\bm{\mu}}W^{(-\bm{t})}_{A}(\bm{\mu})]$. This implies that the complexity of the estimation procedure, determined by the number of samples to achieve a desired precision, depends directly on the range of the function $W^{(-\bm{t})}_{A}$ and the negativity of the PQD $W^{(\bm{t})}_{\rho}$.
@@
Note that in Algorithm 1, $\rho(x)$ is used as a source of samples, while the estimator is mainly built from $\rho(x^{\prime})$ (see Fig. 2). On the other hand, in Algorithm 2, both $U(x)\rho_{\rm{in}}(x)U(x)^{\dagger}$ and $U(x^{\prime})\rho_{\rm{in}}(x^{\prime})U(x^{\prime})^{\dagger}$ are used as sources of samples, while the estimators are mainly built from $\Pi(x)$ and $\Pi(x^{\prime})$ (see Fig. 3). The first method works well for generic cases, while the second method takes the advantage of the specific state preparation. We derive generic conditions for the efficiency of both methods in the next section, and we give concrete applications of these algorithms in Section V.
IV.2 Conditions for efficiency
The efficiency of the previous methods is based on assuming that all the PQDs and their negativity can be efficiently computed, and classical efficient sampling from the probability density is possible. However, if efficient computation of the PQDs of the data encoding states is not possible, one may express the PQDs in terms of the initial product states and the transition function describing the encoding circuit, as we show in Appendix A.
We now present sufficient conditions for the efficiency of the above estimation procedures. Following either method, we obtain a classical estimate of the kernel function $K(x,x^{\prime})$ with additive error $\epsilon=1/\mathrm{poly}(m)$, with an exponentially small probability of failure, using $N=\mathrm{poly}(m)$ samples, provided that both the negativity $\mathcal{N}\big{(}W^{(\bm{t})}_{\rho}\big{)}$ and the range $\mathcal{R}(W^{(-\bm{t})}_{A})$, defining the range of the estimator $\mathcal{R}(E)$ given by Eq. (21), are polynomial in $m$. Thus, to check this condition, one should look for ordering parameters such that $\mathcal{R}(E)$ is minimized. Notice that having non-negative PQD does not necessarily imply the efficient classical estimation of the kernel function, since the other contributing range factor must also be considered in the minimization.
@@ IV.2 Conditions for efficiency
where |$\ket{n,\mu}=D(\mu)\ket{n}$ are displaced number states. If $t\geq 0$ we have $(t-1)/(t+1)\leq 1$. As the operator norm $\|A_{1}\|\leq 1$, the maximum and the minimum values of the single-mode ($-t$)-PQD are determined by the largest and the smallest eigenvalues in Eq. (22):
@@
Therefore, for $t=0$ the values of the Wigner function are between $-2/\pi$ and $2/\pi$, and as $t$ decreases the length of the interval gets smaller, down to $1/\pi$ for the Husimi function when $t=1$. Notice that for $t<0$ the operator $\Delta^{(-t)}(\mu)$ has infinite eigenvalues and hence values of the ($-t$)-PQD are not bounded in general. Generalizing Eq. (23) to the $m$-mode case, we have $\Delta^{(-\bm{t})}(\bm{\mu})=\otimes_{j=1}^{m}\Delta^{(-t_{j})}(\mu_{j})$, and for $t_{j}\geq 0$,
$\bigg{(}\!\frac{t_{\text{min}}-1}{t_{\text{min}}+1}\!\bigg{)}\!\prod_{j=1}^{m}\!\frac{2}{\pi(1+t_{j})}\leq W^{(-\bm{t})}_{A}(\bm{\gamma})\leq\!\prod_{j=1}^{m}\!\frac{2}{\pi(1+t_{j})},$ (24)
@@
V Examples
In this section, we apply our previous findings to various examples. Firstly, we illustrate the use of Algorithm  1 in Section V.1 for estimating quantum kernels based on output states of linear optical networks. Secondly, in Section V.2, we show that there are cases where direct classical computation of the quantum kernel is efficient, such as for kernels based on Gaussian states. Finally, we illustrate the use of Algorithm 2 in Section V.3 for a large class of quantum kernels based on non-Gaussian states obtained through the measurement of a subset of modes of Gaussian states, including adaptive measurement strategies.
V.1 Output states of linear optical networks

### v2

@@ Abstract
Quantum kernel methods are a proposal for achieving quantum computational advantage in machine learning. They are based on a hybrid classical-quantum computation where a function called the quantum kernel is estimated by a quantum device while the rest of computation is performed classically. Quantum advantages may be achieved through this method only if the quantum kernel function cannot be estimated efficiently on a classical computer. In this paper, we provide sufficient conditions for the efficient classical estimation of quantum kernel functions for bosonic systems. These conditions are based on phase-space properties of data-encoding quantum states associated with the quantum kernels: negative volume, non-classical depth, and excess range, which are shown to be three signatures of phase-space negativity. We consider quantum optical examples involving linear-optical networks with and without adaptive non-Gaussian measurements, and investigate the effects of loss on the efficiency of the classical simulation. Our results underpin the role of the negativity in phase-space quasi-probability distributions as an essential resource in quantum machine learning based on kernel methods.
1 Introduction
@@
One approach to address this challenge, especially for quantum sampling problems, has been based on the use of quasi-probability distributions [11, 12, 13, 14, 15, 16, 17]. In this approach, a classical description of a quantum experiment is possible as long as one can find probabilistic descriptions for the input state, the evolution, and the measurement in terms of non-negative quasi-probability distributions. Also, the relation between the negative volume of quasi-probability distributions and the computational overhead in estimating an output probability of a quantum circuit using those distributions has been investigated [16, 17].
@@
In this paper, we build upon these results to quantify the sampling complexity in estimating quantum kernels in terms of the negativity of the associated quasi-probability distributions. Quantum kernel methods have been proposed as a promising near-term application of quantum computers [18], and the non-classicality of quantum kernels has been previously investigated in [19], based on phase-space inequalities from [20]. Hereafter, we go beyond these conceptual characterizations and present sufficient conditions for efficient classical estimation of kernel functions in continuous-variable quantum machine learning. Informally, we show that one can achieve the same performance as a quantum circuit by using classical Monte Carlo-type sampling techniques if the negativity of the phase-space quasi-probability distributions (PQD) of data-encoding quantum states is limited. This condition is independent of the measurement strategy and holds even if sampling from the output probability distribution of the circuit is classically hard, such as in boson sampling problems.
@@
In light of this, we identify the negativity of PQDs as a necessary resource to achieve quantum computational speedups in machine learning based on kernel methods. More precisely, we obtain classical algorithms with fine-grained sample complexity based on specific phase-space properties, namely negative volume [21, 22], non-classical depth [23, 24] and excess ranges. While the negative volume is a direct measure of the negativity of a PQD, the non-classical depth indicates the amount of thermal noise necessary to render a PQD non-negative, and we further show that excess range of PQDs can also be understood a signature of phase-space negativity. As illustrating examples, we discuss efficient classical estimation of quantum kernel functions that are based on Gaussian states, non-Gaussian output states of linear optical networks, and partially measured Gaussian states, including adaptive measurement strategies.
@@
Our results involve a new approach to classical estimation of quantum kernel functions beyond probability estimation [16, 17] that takes into account the structure of state preparation. A surprising consequence is that estimating quantum kernel functions based on a large class of partially measured Gaussian states can be done efficiently by classical computers despite these states being highly non-Gaussian (see Theorem 1).
@@
The structure of the paper is as follows: in Section 2, we set the stage and provide some background on quantum kernel methods; in Section 3, we derive general expressions for quantum kernel functions for bosonic systems based on PQDs (see Lemma 1); in Section 4, we introduce classical Monte Carlo sampling methods and derive two phase-space-inspired classical algorithms for estimating quantum kernel functions (see Algorithms 1 and 2), together with rigorous performance guarantees; we illustrate the flexibility and applicability of our algorithms in Section 5, by applying them to several examples of high significance in bosonic quantum information processing, identifying the relevant quantum computational resources in each setting; we conclude in Section 6.
2 Quantum Kernel Methods
@@
3 Kernel functions in phase space
@@
where $\bm{\alpha}$ is a row vector of $m$ complex numbers, where $\text{d}^{2m}\!\bm{\bm{\alpha}}\,=\mathrm{d}^{m}{\,\Re(\bm{\alpha})}\mathrm{d}^{m}{\,\Im(\bm{\alpha})}$, and where the $\bm{s}$-ordered phase-space quasi-probability distributions [($\bm{s}$)-PQDs] are defined by
@@
Here, $\bm{\alpha},\bm{\xi}$ are row vectors of $m$ complex numbers, the diagonal matrix $\bm{s}=\mathrm{diag}(s_{1},s_{2},\ldots,s_{m})$ specifies the ordering for each mode, and $D(\bm{\xi})=\exp(\bm{\xi}\bm{a}^{\dagger}-\bm{a}\bm{\xi}^{\dagger})$ is the $m$-mode displacement operator with $\bm{a}=(a_{1},\ldots,a_{m})$ and $\bm{a}^{\dagger}=(a_{1}^{\dagger},\ldots,a_{m}^{\dagger})^{T}$ being vectors of annihilation and creation operators, respectively. Note that these formulas are formally valid for more general ordering matrices, but we restrict to diagonal ones for simplicity.
@@
$\mathrm{Tr}[\rho_{\Pi(x)}\rho_{\Pi(x^{\prime})}]$ $=\pi^{m+2k}\!\!\int_{\bm{\alpha},\bm{\beta}\in\mathbb{C}^{k}}\!\!\!\!\!\!\!\!\!\!\mathrm{d}^{2k}\bm{\alpha}\,\mathrm{d}^{2k}\bm{\beta}\,W_{\Pi(x)}^{(-\bm{u})}(\bm{\alpha})W_{\Pi(x^{\prime})}^{(\bm{v})}(\bm{\beta})$ (13)
×$\quad\quad\quad\times\int_{\bm{\gamma}\in\mathbb{C}^{m}}\!\!\!\!\!\!\!\mathrm{d}^{2m}\bm{\gamma}\,W_{U(x)\rho_{\rm{in}}(x)U(x)^{\dagger}}^{(\bm{u}\oplus\bm{s})}(\bm{\alpha},\bm{\gamma})W_{U(x^{\prime})\rho_{\rm{in}}(x^{\prime})U(x^{\prime})^{\dagger}}^{(-\bm{v}\oplus-\bm{s})}(\bm{\beta},\bm{\gamma}).$
@@
$=\mathrm{Tr}[\sigma(\Pi\otimes\Delta^{(\bm{s})}(\bm{\gamma}))]$
$=\pi^{k}\!\!\int_{\bm{\alpha}\in\mathbb{C}^{k}}\!\!\!\!\!\mathrm{d}^{2k}\bm{\alpha}W_{\Pi}^{(-\bm{u})}\!(\bm{\alpha})$ (15)
×$\quad\quad\quad\times\mathrm{Tr}[\sigma(\Delta^{(\bm{u})}\!(\bm{\alpha})\!\otimes\!\Delta^{(\bm{s})}\!(\bm{\gamma}))]$
@@
4 Classical estimation of kernel functions
In this section, we describe two classical algorithms using Monte Carlo sampling methods to estimate kernel functions based on the expressions obtained in Lemma 1. The first Algorithm 1 is obtained by treating kernel estimation as a probability estimation task, and following the approach of [17, 16, 32] (see Fig. 2). The second Algorithm 2 is a non-trivial generalisation which takes advantage of the specifics of state preparation, and uses the first algorithm as a subroutine (see Fig. 3).
4.1 Monte Carlo estimation
@@
where $\mathcal{N}(W^{(\bm{t})}_{\rho}):=\int_{\mathbb{C}^{n}}\text{d}^{2n}\!\bm{\mu}\,\,\big{|}W^{(\bm{t})}_{\rho}(\bm{\mu})\big{|}$ is the negative volume, a measure of negativity in the ($\bm{t}$)-PQD. We note that $\mathcal{N}(W^{(\bm{t})}_{\rho})=1$ if the distribution is non-negative. Next, using the quasi-probability distribution $W^{(-\bm{t})}_{A}(\bm{\mu})$ representing operator $A$, we introduce the estimator
@@
then with probability at least $1-\delta$ the classical estimate of the expectation value has additive error less than $\epsilon$. Here, the range of the estimator is bounded as
$\mathcal{R}(E)\leq 2\mathcal{N}\big{(}W^{(\bm{t})}_{\rho}\big{)}\mathcal{R}(W^{(-\bm{t})}_{A}),$ (21)
with $\mathcal{R}(W^{(-\bm{t})}_{A})=\pi^{n}[\max_{\bm{\mu}}W^{(-\bm{t})}_{A}(\bm{\mu})-\min_{\bm{\mu}}W^{(-\bm{t})}_{A}(\bm{\mu})]$. This implies that the complexity of the estimation procedure, determined by the number of samples to achieve a desired precision, depends directly on the range of the function $W^{(-\bm{t})}_{A}$ and the negative volume of the PQD $W^{(\bm{t})}_{\rho}$. Note that $\pi^{n}\max_{\bm{\mu}}|W^{(-\bm{t})}_{A}(\bm{\mu})|\leq\mathcal{R}(W^{(-\bm{t})}_{A})\leq 2\pi^{n}\max_{\bm{\mu}}|W^{(-\bm{t})}_{A}(\bm{\mu})|$, so we can alternatively use the extremal values rather than the range without affecting the scaling of the sample complexity.
@@
Note that in Algorithm 1, $\rho(x)$ is used as a source of samples, while the estimator is mainly built from $\rho(x^{\prime})$ (see Fig. 2). On the other hand, in Algorithm 2, both $U(x)\rho_{\rm{in}}(x)U(x)^{\dagger}$ and $U(x^{\prime})\rho_{\rm{in}}(x^{\prime})U(x^{\prime})^{\dagger}$ are used as sources of samples, while the estimators are mainly built from $\Pi(x)$ and $\Pi(x^{\prime})$ (see Fig. 3). The first method works well for generic cases, while the second method takes the advantage of the specific state preparation. We derive generic conditions for the efficiency of both methods in the next section, and we give concrete applications of these algorithms in Section 5.
4.2 Conditions for efficiency
The efficiency of the previous methods is based on assuming that all the PQDs and their negative volume can be efficiently computed, and classical efficient sampling from the probability density is possible. However, if efficient computation of the PQDs of the data encoding states is not possible, one may express the PQDs in terms of the initial product states and the transition function describing the encoding circuit, as we show in Appendix A.
We now present sufficient conditions for the efficiency of the above estimation procedures. Following either method, we obtain a classical estimate of the kernel function $K(x,x^{\prime})$ with additive error $\epsilon=1/\mathrm{poly}(m)$, with an exponentially small probability of failure, using $N=\mathrm{poly}(m)$ samples, provided that both the negative volume $\mathcal{N}\big{(}W^{(\bm{t})}_{\rho}\big{)}$ and the range $\mathcal{R}(W^{(-\bm{t})}_{A})$, bounding the range of the estimator $\mathcal{R}(E)$ as in Eq. (21), are polynomial in $m$. Thus, to check this condition, one should look for ordering parameters such that $\mathcal{R}(E)$ is minimized.
Notice that having a nonnegative PQD does not necessarily imply efficient classical estimation of the kernel function, since the other contributing range factor must also be considered in the minimization. In fact, the range of PQD is also related to the negativity involved in kernel estimation experimental procedures. As $A$ is a bounded positive operator, it can be viewed as a POVM element of a two-outcome measurement $\{A,\bar{A}=I-A\}$. Therefore, estimation of $\mathrm{Tr}[\rho\bar{A}]$ results in estimation of the complement result of the measurement, $\mathrm{Tr}[\rho A]=1-\mathrm{Tr}[\rho\bar{A}]$. However, since $W^{(-\bm{t})}_{\bar{A}}(\bm{\mu})=1/\pi^{m}-W^{(-\bm{t})}_{A}(\bm{\mu})$, we can see that if the range of $W^{(-\bm{t})}_{A}(\bm{\mu})$ is greater than $1/\pi^{m}$, then either it takes negative values, or the PQD $W^{(-\bm{t})}_{\bar{A}}(\bm{\mu})$ involved in the estimation of $\mathrm{Tr}[\rho\bar{A}]$ becomes negative. In other terms, the excess range of a PQD is a signature of phase-space negativity contributing to the sample complexity of our classical algorithms.
@@ 4.2 Conditions for efficiency
where |$\ket{n,\mu}=D(\mu)\ket{n}$ are displaced number states. If $t\geq 0$, we have $(t-1)/(t+1)\leq 1$. As the operator norm $\|A_{1}\|\leq 1$, the maximum and the minimum values of the single-mode ($-t$)-PQD are determined by the largest and the smallest eigenvalues in Eq. (22):
@@
Therefore, for $t=0$ the values of the Wigner function are between $-2/\pi$ and $2/\pi$, and as $t$ decreases the length of the interval becomes smaller, down to $1/\pi$ for the Husimi function when $t=1$. Notice that for $t<0$ the operator $\Delta^{(-t)}(\mu)$ has infinite eigenvalues, and therefore the values of ($-t$)-PQDare not bounded in general.
Generalizing Eq. (23) to the $m$-mode case, we have $\Delta^{(-\bm{t})}(\bm{\mu})=\otimes_{j=1}^{m}\Delta^{(-t_{j})}(\mu_{j})$, and for $t_{j}\geq 0$,
$\bigg{(}\!\frac{t_{\text{min}}-1}{t_{\text{min}}+1}\!\bigg{)}\!\prod_{j=1}^{m}\!\frac{2}{\pi(1+t_{j})}\leq W^{(-\bm{t})}_{A}(\bm{\gamma})\leq\!\prod_{j=1}^{m}\!\frac{2}{\pi(1+t_{j})},$
@@
5 Examples
In this section, we apply our previous findings to various examples. Firstly, we illustrate the use of Algorithm  1 in Section 5.1 for estimating quantum kernels based on output states of linear optical networks. Secondly, in Section 5.2, we show that there are cases where direct classical computation of the quantum kernel is efficient, such as for kernels based on Gaussian states. Finally, we illustrate the use of Algorithm 2 in Section 5.3 for a large class of quantum kernels based on non-Gaussian states obtained through the measurement of a subset of modes of Gaussian states, including adaptive measurement strategies.
5.1 Output states of linear optical networks